On January 1, 2021, the Mason Manufacturing Company began construction of a building to be used as its office headquarters. The building was completed on September 30, 2022. Expenditures on the project were as follows:

January 1, 2021 $ 1,740,000
March 1, 2021 1,380,000
June 30, 2021 1,580,000
October 1, 2021 1,380,000
January 31, 2022 387,000
April 30, 2022 720,000
August 31, 2022 1,017,000


On January 1, 2021, the company obtained a $4,300,000 construction loan with a 12% interest rate. The loan was outstanding all of 2021 and 2022. The company’s other interest-bearing debt included two long-term notes of $3,000,000 and $7,000,000 with interest rates of 8% and 10%, respectively. Both notes were outstanding during all of 2021 and 2022. Interest is paid annually on all debt. The company’s fiscal year-end is December 31.

Required:
1. Calculate the amount of interest that Mason should capitalize in 2021 and 2022 using the specific interest method.
2. What is the total cost of the building?
3. Calculate the amount of interest expense that will appear in the 2021 and 2022 income statements.
Answer #1
2021 2022
Interest Capitalized $               4,83,000 $         6,06,960
Interest Expense $               9,73,000 $         8,49,040
Cost of Building = $       92,93,960
Workings:
Expenditure for 2021
Jan 1,2021 $            17,40,000 X 12/12 = $        17,40,000
March 1, 2021 $            13,80,000 X 10/12 = $        11,50,000
June 30, 2021 $            15,80,000 X 6/12 = $          7,90,000
October 1, 2021 $            13,80,000 X 3/12 = $          3,45,000
$            60,80,000 $        40,25,000
Interest Capitalized in 2021
$            40,25,000 X 12.00% = $          4,83,000
Expenditure for 2022
Jan 1, 2022 ($680000+$483000) $            65,63,000 X 9/9 = $        65,63,000
Jan 31, 2022 $               3,87,000 X 8/9 = $          3,44,000
April 30, 2022 $               7,20,000 X 5/9 = $          4,00,000
August 31. 2022 $            10,17,000 X 1/9 = $          1,13,000
$            86,87,000 $        74,20,000
Interest Capitalized in 2020
$                       74,20,000
Less: $                       43,00,000 X 12.00% X 9/12 = $          3,87,000
$                       31,20,000 X 9.40% X 9/12 = $          2,19,960
Interest Capitalized in 2020 = $          6,06,960
Weighted Average rate of all debt:-
$            30,00,000 X 8% = $          2,40,000
$            70,00,000 X 10% = $          7,00,000
$         1,00,00,000 $          9,40,000
Weighted Average rate of all debt = 9.40%
($940000/ 100,00,000)
Interest Expense for 2021:
Total Interest Incurred ($4300000*12%)+($3000000*8%)+($7000000*10%) = $        14,56,000
Less : Interest Capitalized = $          4,83,000
2021 Expense = $          9,73,000
Interest Expense for 2022:
Total Interest Incurred = $        14,56,000
Less : Interest Capitalized = $          6,06,960
2022 Expense = $          8,49,040
Cost of Building
Expenditure for 2021 = $        60,80,000
Interest Capitalized in 2021 = $          4,83,000
Expenditure for 2022 = $        21,24,000
Interest Capitalized in 2022 = $          6,06,960
Cost of Building = $        92,93,960